OpenStudy (anonymous):

A batter hits a softball straight up at 62 m/s how high up does it go?

OpenStudy (anonymous):

physics problem: Initial velocity is 62 m/s and we want to find the max height. Final velocity at this height will be 0 m/s. The acceleration will be the negative force of gravity. Can you solve it now?

OpenStudy (anonymous):

no...-9.8 is the force of accelleratiion due to gravity...not sure of the formula

OpenStudy (anonymous):

where x(f)-x(0) = change is x or distance

OpenStudy (anonymous):

v(f)^2=v(0)^2+2a(d)

OpenStudy (anonymous):

uhhh..k...still not sure...anyone else?

OpenStudy (anonymous):

Hopefully somebody can collaborate my formula given, but that is what I remember it to be. You know everything but d to solve the equation. Good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

v² = u² + 2as where v = final velocity u = initial velocity s = displacement a = acceleration (in this case due to gravity) Initially, velocity = 62m/s and finally at maximum height, velocity = 0 m/s v² = u² + 2as where v = 0 m/s u = 62 m/s s = ? a = (gravity) -9.81 Replacing values : 0 = 62² + (2 x s x -9.81) 62² = 2s x (9.81) s = 62² / (2 x 9.81) = 195.922528 m